### v1

# FedMLLM: Federated Fine-tuning MLLM on Multimodal Heterogeneity Data

Abstract
Multimodal Large Language Models (MLLMs) have made significant advancements, demonstrating powerful capabilities in processing and understanding multimodal data. Fine-tuning MLLMs with Federated Learning (FL) allows for expanding the training data scope by including private data sources, thereby enhancing their practical applicability in privacy-sensitive domains. However, current research remains in the early stage, particularly in addressing the multimodal heterogeneities in real-world applications. In this paper, we introduce a benchmark for evaluating various downstream tasks in the federated fine-tuning of MLLMs within multimodal heterogeneous scenarios, laying the groundwork for the research in the field. Our benchmark encompasses two datasets, five comparison baselines, and four multimodal scenarios, incorporating over ten types of modal heterogeneities. To address the challenges posed by modal heterogeneity, we develop a general FedMLLM framework that integrates four representative FL methods alongside two modality-agnostic strategies. Extensive experimental results show that our proposed FL paradigm improves the performance of MLLMs by broadening the range of training data and mitigating multimodal heterogeneity. Code is available at https://github.com/1xbq1/FedMLLM
1 Introduction
Multimodal Large Language Models (MLLMs), which regard Large Language Models (LLMs) as core components for performing various multimodal tasks, have rapidly emerged as a research hotspot [39]. Their exceptional performance and potential as a pathway to artificial general intelligence contribute to this trend. However, the high computational costs and the requirement for large amounts of training data are significant challenges that hinder the further development and application of MLLMs in the real world [37, 24]. For the former, various lightweight MLLMs [37, 23] and Parameter Efficient Fine-Tuning (PEFT) [11, 20] methods, such as Low-Rank Adaptation (LoRA) [11], have been investigated. For the latter, a practical solution is to extend the data scope from public datasets to unused private data via Federated Learning (FL) [27, 15].
However, there is limited research on fine-tuning MLLMs under the FL framework [43], with most efforts focused on integrating LLMs with FL [38, 19, 42]. Compared to LLMsFL, MLLMsFL supports more modal inputs, making it more applicable to real-world scenarios [39]. For example, in natural disasters or emergency situations (e.g., earthquakes, floods), it is essential to integrate multi-modal data such as images or videos (on-site photos), text (social media reports), and audio (distress signals). MLLMs+FL enable distributed collaboration across organizations while safeguarding data privacy, offering a more comprehensive understanding of disasters through multi-modal modeling. In contrast, LLMs+FL are limited to text, missing key information from images and audio. However, the multimodal data, as the input of FL framework of multi-client collaborative training, also brings about more complex modal heterogeneity. As shown in Fig. 1, multimodal scenarios can be categorized as the aligned modality [9], and the non-aligned modality (i.e., missing modality [26, 3], cross modality [36], and hybrid modality [40, 5]). Only the heterogeneity of the aligned modality is consistent with that of a single modality, while the non-aligned multimodal situations exhibit additional modal heterogeneity.
To eliminate the bias caused by modal heterogeneity, we design two simple yet effective modal-agnostic strategies based on existing FL methods, as illustrated in Fig. 2. First, the prompt strategy instructs LLMs to perform tasks without considering modality, e.g., “…, without considering the modality.”, which considers the effect of instructions on LLMs at the data level [33]. Second, the regularization strategy constrains the update direction. As the client moves from aligned to missing modalities, eventually to extreme single-modality cases, the risk of overfitting to a specific modality gradually increases. To address this, the adaptive regularization strategy adjusts consistency with global weights based on the severity of missing modalities, with stronger consistency for more missing modalities. Meanwhile, only the middle layers are regularized, preserving the ability of the low and high layers to extract modal features and provide task-specific representations.
Combined with the above modal-agnostic strategies, we comprehensively explore the pipeline of federated fine-tuning of MLLM on decentralized multimodal downstream data. Unlike [43], which applies MLLM solely to assist FL in alleviating traditional data heterogeneity and long-tail distribution problems, we focus more on addressing the issue of diverse modal heterogeneity in multimodal scenarios that are closely related to MLLMs. As shown in Fig. 2, the proposed pipeline supports four multimodal scenarios: i. Aligned Modal Scenario refers to all samples of each client existing in an aligned multimodal form. ii. Missing Modal Scenario occurs when a certain modality is absent in some samples of each client. iii. Cross Modal Scenario involves each client having only one modality, while all clients together cover multiple modalities. iv. Hybrid Modal Scenario refers to some clients containing only one modality, while others contain a combination of multiple modalities [9]. In total, we construct over ten cases of modal heterogeneity across two datasets, based on varying heterogeneity degrees and the proportions of non-aligned multi-modalities. Finally, the application of FL in MLLMs significantly improves performance compared to local training, as shown in the four modal results of Fig. 2. Additionally, the two modality-agnostic strategies proposed for the modal heterogeneity problem also contribute positively.
Overall, our contributions can be summarized in three folds. 1) We explore the fine-tuning MLLM benchmark on decentralized multimodal data to better adapt to various downstream tasks in comprehensive multimodal federated scenarios, which highlights a potential development direction for MLLMs. 2) We build a general FedMLLM framework to better address the challenges of multimodal heterogeneity by elegantly incorporating two modality-agnostic strategies on top of multiple FL baselines. 3) We conduct extensive experiments, related to 2 downstream datasets, 5 comparison baselines, and 4 multimodal scenarios, to show that FL paradigm enhances MLLM performance by broadening the training data scope without bells and whistles.
2 Related Work
2.1 Multimodal Large Language Models
The rapid development of Large Language Models (LLMs) has significantly driven the research surge in Multimodal Large Language Models (MLLMs) [39]. Notably, the emergence of commercial models such as GPT-4 [28] and Gemini [32] has attracted widespread attention. Simultaneously, increasingly powerful open-source models [21, 7, 44] have also emerged. BLIP-2 [21] is well-known for its Q-former module, with promising VQA results. Building on BLIP-2 and Vicuna [7], MiniGPT-4 [44], derived from GPT-4 [28], is proposed for image-text pairs. Unlike previous methods, LLaVA [24] is trained on multimodal instruction data generated by GPT-4 to enhance its zero-shot capability.
However, the large number of parameters in MLLMs leads to high computational costs for training and deployment, limiting their broader application. Recently, efficient MLLMs have also gained significant attention within the academia [14], with representative models including Phi [13], Gemma [2], MobileLLM [25], and MiniCPM-V [37]. Compared to previous approaches, MiniCPM-V [37] optimizes architecture, training, and inference to achieve superior performance with fewer parameters. Consequently, we utilize MiniCPM-V as the MLLM for both the client and server in the federated learning process.
2.2 Federated Learning
As a classic framework for privacy protection, Federated Learning (FL) typically allows multiple parties to co-train while keeping the data local [12]. The original FedAvg [27] method aggregates model weights from different clients solely based on sample proportions, which makes it challenging to effectively address complex data heterogeneity. As a result, various variants have emerged: some constrain the update direction of the local model on the client side [22, 16], while others optimize the aggregation method on the server side [30, 10, 31].
Our benchmark includes four representative FL methods for server-side aggregation. FedYogi, FedAdam, and FedAdagrad [30] apply the centralized Yogi [41], Adam [18], and Adagrad [8] optimizers to the server-side aggregation in the FL framework, while FedAvgm [10] uses momentum updates on the server, similar to momentum applied on top of SGD. Because the original FedAvg [27] and FedProx [22] with a global regularization term face training instability during federated fine-tuning of MLLM, the conventional average aggregation method does not converge effectively. As a result, these two methods are not included separately in the comparison baselines.
2.3 Multimodal Federated Learning
Unlike the more established single-modal FL, Multimodal Federated Learning (MFL) has significant potential for exploration [9]. Existing MFL methods can also be categorized based on the different modal scenarios in our benchmark. In aligned modal scenarios, MFL techniques resemble those of single-modal FL, focusing on client-specific and shared knowledge [4]. Some methods also draw from multimodal learning, especially in feature fusion [9]. In missing modal scenarios, some MFL methods rely on pre-trained generative models to fill the gaps in modality [3]. In cross modal scenarios, some MFL methods employ a two-branch structure to separately constrain the learning of specific and shared modalities [36, 35, 29], while hybrid modal MFL methods utilize various graph techniques to capture complex correlations among clients [5, 6].
Typically, MFL extracts features from raw inputs using pre-trained models before training, but its performance is ultimately limited by the effectiveness of these models in feature extraction [9, 36]. With the rise of MLLMs, their application in MFL has become a trend. However, only one study so far has attempted to combine MLLMs with FL on the server side to address issues of data heterogeneity and long-tail distribution [43]. In this paper, we present the first benchmark for federated fine-tuning of MLLMs, targeting diverse multimodal heterogeneities.
3 Multimodal Scenarios Construction
For a comprehensive simulation of multimodal scenarios in the context of FL, we refer to [9, 3, 36]. As shown in Fig. 2, the simulation is divided into four scenarios: Aligned Modal Scenario, Missing Modal Scenario, Cross Modal Scenario, and Hybrid Modal Scenario.
Aligned Modal Scenario. The dataset is sampled using Dirichlet distribution, similar to the traditional partitioning method [9], with coefficients adjusted to control the distribution of samples among different clients. Specifically, for each label, Dirichlet distribution is used to allocate the number of samples to each client; the smaller the coefficient, the greater the difference in sample sizes between clients, resulting in increased data heterogeneity. Here, each client’s sample contains all modalities.
Missing Modal Scenario. The dataset first assigns samples to different clients according to the partitioning strategy of aligned modal scenarios. Currently, the samples in each client are full-modal. By using a uniform distribution for each modality and setting a missing rate of to determine whether to remove that modality, the situation where the samples in each client have different missing modalities is achieved.
Cross Modal Scenario. The dataset is first allocated according to a partitioning strategy similar to the aligned modal scenario and then divided based on the ratio ({, , }) of image-only to text-only clients. For example, in the ratio, 3 out of 10 clients are randomly assigned the image modality, while the other 7 clients receive the text modality. The other ratios follow the same pattern. This approach allows each client to have a single modality while ensuring that all clients collectively have multiple modalities.
Hybrid Modal Scenario. After applying a partitioning method consistent with the aligned modality scenario, the dataset follows a Bernoulli distribution for each client’s modality retention probability , which is the same for each modality. Once the partitioning is complete, a single client may consist of either a single modality or a combination of multiple modalities, with internal samples maintaining consistency in modality. To avoid an invalid dataset missing both modalities on the client side, in such cases, either the image or text modality is randomly retained with a 50% probability.
As shown in Fig. 3 and Fig. 4, we present visualizations of heterogeneity in the aligned modal scenario on the CrisisMMD dataset and of multimodal heterogeneities across four multimodal scenarios on the Hateful-Memes dataset, respectively. More visualization results can be found in the Supplementary Material. Overall, these four multimodal scenarios initially simulate simple situations with various modal inputs in real life.
4 FedMLLM Framework
4.1 Overview of FedMLLM
As shown in the architecture part of Fig. 2, the proposed FedMLLM is divided into two parts: the cloud (server-side) and the terminal (client-side). Both parts use the same MLLM [37] and share a homogeneous model architecture. In FedMLLM, we address an optimization problem similar to that of traditional FL [27], as follows,
where represents the loss function of client among a total of clients, specifically given by , with denoting the data distribution for the client. Here, the model parameters are given by , where refers to the frozen pretrained weights, and is the trained LoRA component [11] of the LLM.
The FL process of FedMLLM starts on the server side, which sends either the initialized LoRA weights (in the first round) or the aggregated weights (in subsequent rounds) to the selected clients to initialize their LoRA weights. Then, client fine-tunes the initialized LoRA weights on the local dataset , where , , denote the instruction token sequences, multimodal input token sequences, and ground truth response token sequences, respectively, with . The loss function [39] maximizes the probability of the current input sequence generating the next correct token, as follows,
where is the length of , and represents the model parameters of client at epoch of round . Once local training is complete, the updated LoRA weights from the selected clients are uploaded to the server for aggregation. At this stage, one training round has been completed, and the FL training process needs to be repeated for several more rounds until the model converges.
4.2 Modality-agnostic Strategies
Prompt Strategy. In decentralized multimodal data scenarios, different modality conditions among clients further complicate the data heterogeneity problem. Lacking modalities may introduce additional modality bias into the model’s learning process, affecting the convergence of the global model. A common solution is to address this issue at the data source by using a large-scale pre-trained model to supplement the modality [3]. Nevertheless, it often incurs additional costs. To avoid these costs, we aim to mitigate the impact of modality changes on task outcomes by explicitly prompting model training. Inspired by this, we augment the original prompt by introducing a modality-agnostic prompt strategy to enhance the model’s robustness and generalization ability, as shown in Tab. 1.
Regularization Strategy. The regularization term originally introduced by FedProx [22] constrains the direction of local updates toward the global model, serving as a general method for mitigating heterogeneity. However, the average aggregation strategies used in FedProx and FedAvg [27] can exacerbate instability during federated fine-tuning of MLLM, leading to a gradual increase in loss and its variance (see in Supplementary Material). In contrast, FedAvgM employs server momentum [10], while FedAdam, FedYogi, and FedAdagrad use adaptive optimizers [30], allowing them to handle instability more effectively.
| Method | ||||||
| Hateful-Memes | CrisisMMD | Hateful-Memes | CrisisMMD | Hateful-Memes | CrisisMMD | |
| Local | 71.72 | 55.56 | 67.38 | 45.78 | 67.71 | 37.67 |
| FedYogi | 74.56 | 61.21 | 71.53 | 61.56 | 71.13 | 60.90 |
| FedAdam | 74.67 | 61.04 | 73.17 | 59.52 | 71.74 | 61.01 |
| FedAvgM | 74.39 | 56.69 | 74.55 | 56.76 | 62.51 | 32.09 |
| FedAdagrad | 73.90 | 60.88 | 71.99 | 60.66 | 72.69 | 59.61 |
| Method | ||||||
| Hateful-Memes | CrisisMMD | Hateful-Memes | CrisisMMD | Hateful-Memes | CrisisMMD | |
| Local | 66.83 | 33.34 | 66.94 | 31.88 | 66.90 | 31.56 |
| FedYogi | 69.96 | 60.06 | 69.12 | 59.16 | 74.17 | 54.57 |
| FedAdam | 72.52 | 60.61 | 72.35 | 57.96 | 73.15 | 54.33 |
| FedAvgM | 63.07 | 33.17 | 59.92 | 34.57 | 58.69 | 32.99 |
| FedAdagrad | 72.67 | 56.65 | 71.05 | 56.30 | 70.94 | 56.17 |
| Method | ||||||
| Hateful-Memes | CrisisMMD | Hateful-Memes | CrisisMMD | Hateful-Memes | CrisisMMD | |
| Local | 67.04 | 25.92 | 66.38 | 28.82 | 67.38 | 31.42 |
| FedYogi | 67.02 | 47.46 | 75.09 | 58.03 | 76.02 | 60.97 |
| FedAdam | 67.35 | 51.88 | 74.78 | 58.20 | 74.56 | 60.91 |
| FedAvgM | 61.50 | 31.31 | 63.91 | 38.67 | 58.77 | 33.41 |
| FedAdagrad | 69.33 | 49.22 | 72.47 | 51.21 | 72.35 | 51.55 |
| Method | ||||||
| Hateful-Memes | CrisisMMD | Hateful-Memes | CrisisMMD | Hateful-Memes | CrisisMMD | |
| Local | 67.10 | 35.37 | 67.66 | 33.52 | 67.68 | 33.70 |
| FedYogi | 66.36 | 58.60 | 69.60 | 60.82 | 70.02 | 58.57 |
| FedAdam | 70.43 | 60.84 | 70.68 | 61.91 | 71.26 | 59.39 |
| FedAvgM | 60.43 | 47.92 | 62.12 | 43.27 | 60.35 | 39.07 |
| FedAdagrad | 69.69 | 57.36 | 72.18 | 57.87 | 69.93 | 53.45 |
To reduce the negative effect of modal heterogeneity, a straightforward approach is to integrate the FedProx regularization term into the stable FL baselines. This traditional regularization strategy with a fixed coefficient is not suitable for various non-aligned modal scenarios. Specifically, as the issue of missing modalities becomes more severe in such scenarios, it is more reasonable to strengthen the regularization to further prevent overfitting to a particular modality. At the same time, to preserve both the low-level modality-specific and high-level task-specific information, we only regularize the intermediate modality-agnostic layers. Therefore, an adaptive regularization strategy is presented, as follows,
where is an adaptive coefficient, is the layer-wise mask, is the transpose operation, is a layer-count hyperparameter, and represents the server-side aggregated LoRA weight. In cross modal scenario, suppose is set to . When the missing rate is in missing modal scenario, is linearly adjusted to to adaptively adjust the strength of agreement with the global model. The hybrid modal scenario combines aligned, missing, and cross modal scenarios. Thus, the setting follows these scenarios, with no regularization in the aligned case. In practice, is calculated from the client’s private dataset.
5 Experiments
5.1 Experimental Settings
Datasets. To stay updated on the latest trends, modern society is flooded daily with social media information from various regions and across diverse scopes. However, effectively identifying and reducing misinformation and harmful content within these platforms has become a critical task in current research [9]. Moreover, social media information from different regions and privacy levels has driven research on FL in this field. Therefore, in evaluating the FedMLLM framework, we utilize two multimodal social media datasets, Hateful-Memes [17] and CrisisMMD [1], containing image-text pairs.
Hateful-Memes [17] contains a total of 10k memes focused on detecting hate speech in multimodal memes, with the training set, dev set, and test set comprising 85%, 5%, and 10%, respectively. In this study, we only use the training set and the test set. The dataset includes a total of five types of data: multimodal hate, unimodal hate, benign image, benign text, and random not-hateful data. The task aims to determine if the given image and text are hateful, a binary classification problem. The evaluation metric used is the Area Under the Receiver Operating Characteristic Curve (ROC AUC), as recommended by the authors.
CrisisMMD [1] includes 18.1k tweets from major natural disasters that occurred in 2017 across different parts of the world. We use a training set of 13.6k tweets and a test set of 2.2k tweets, excluding the dev set. The dataset comprises eight humanitarian categories for classification: affected individuals, infrastructure and utility damage, injured or dead people, missing or found people, rescue volunteering or donation effort, vehicle damage, other relevant information, not humanitarian. The evaluation metric is the F1 score [9].
Configurations. For the model configuration, the efficient MiniCPM-V-2_6-int4 [37] is adopted, with Qwen2 [34] as the LLM, and LoRA parameters are set to a rank of 8, alpha of 8, and dropout of 0.05. The maximum sequence length for the model is 1200. Formatting instructions can be found in the Supplementary Material. In the FL settings, there are 50 training rounds with 10 clients, sampling 2 clients per round, each training for one epoch. The initial learning rate is set to 2e-5 and adjusted using a cosine scheduler with a warmup ratio of 0.01. In the local baseline, the number of training epochs is 5. The data split method is described in Section Sec. 3. In the non-aligned modal scenario, partitioning first uses from the aligned modal scenario, followed by the specific non-aligned modal partition method. All experiments use a single NVIDIA A40 or A100 GPU with 617M trainable parameters out of 8716M total, using around 30GB of GPU memory.
5.2 Performance Evaluation
Comparison of FL Baselines and Local. Tab. 2-Tab. 5 sequentially present the results of five comparison baselines across two downstream datasets under 12 different levels of heterogeneity in the aligned modal, missing modal, cross modal, and hybrid modal scenarios. Compared to the “Local” baseline that averages results from clients independently training on local data, the four FL baselines perform better in heterogeneous conditions, especially FedYogi [30], FedAdam [30], and FedAdagrad [30] that utilize adaptive optimizers. This indicates that collaborative training in FL, which indirectly expands the range of training data, further improves MLLM performance in heterogeneous scenarios.
Non-Aligned vs. Aligned Modal Scenarios. Fig. 5 illustrates how the performance of various FL baselines shifts from the aligned modal scenario () to non-aligned modal scenarios. In the multi-class CrisisMMD dataset, the performance of FedYogi, FedAdam, and FedAdagrad is generally lower than that in aligned modal scenarios. For example, the performance of FedYogi decreases by 6.33% when heterogeneity is compared to . However, in the binary-class Hateful-Memes dataset, FedYogi and FedAdam even show performance improvements under missing and cross modal scenarios. For instance, FedAdam improves by 3.04% when heterogeneity is I-5:T-5 compared to . On the contrary, FedAvgM [10] is more affected by data heterogeneity at than at , as shown in Tab. 2. Overall, in binary-class classification tasks, moderate modal heterogeneity functions similarly to data augmentation, enhancing model robustness to some extent and thereby improving performance. However, in multi-class classification tasks, modal heterogeneity primarily has a negative impact, leading to a decline in the performance of FL methods.
Influence of Different Multimodal Scenarios. Fig. 6 shows the overall performance differences across different modal scenarios and reveals the influence of these scenarios on performance. We can see that: 1) Each modal scenario contains certain heterogeneous conditions that significantly degrade the performance of FedMLLM on both datasets. 2) Hybrid modal scenario outperforms other non-aligned modal scenarios in overall performance on the CrisisMMD dataset. However, on the Hateful-Memes dataset, its best performance is generally lower than that of other non-aligned modal scenarios. 3) The performance distributions of the Missing and Hybrid modal scenarios exhibit similar trends, whereas the Cross modal scenario has the widest performance distribution, encompassing both the best performance on the Hateful-Memes dataset and the worst performance on the CrisisMMD dataset.
| Scenario | Hateful-Memes | CrisisMMD |
| Missing | 70.94 | 56.17 |
| 71.66(+0.72) | 57.20(+1.03) | |
| Cross | 69.33 | 49.22 |
| 71.17(+1.84) | 49.56(+0.34) | |
| Hybrid | 69.69 | 53.45 |
| 70.59(+0.90) | 58.46(+5.01) |
Results on Modality-agnostic Strategies. Tab. 6 and Tab. 7 demonstrate the performance improvement of two modality-agnostic strategies across various non-aligned modal scenarios and FL methods, respectively. The listed non-aligned scenarios are those in which performance is most negatively affected, while the addition of both modality-agnostic strategies improves performance in these scenarios. For different FL methods, both strategies also help mitigate the negative effects of modal heterogeneity. Specifically, FedAdagrad, which is solely negatively affected by modal heterogeneity, improves with both strategies, showing a 5.01% gain on the CrisisMMD dataset in the Hybrid modal scenario. Furthermore, Tab. 8 presents ablation studies showing that prompt and regularization strategies effectively mitigate modality heterogeneity, yielding the best performance when applied together. Fig. 7 shows the ablation study for the and layers in adaptive regularization, with final values of and set to 0.1 and 4, respectively.
| Method | Hateful-Memes | CrisisMMD |
| FedYogi | 67.02 | 47.46 |
| 69.80(+2.78) | 49.46(+2.00) | |
| FedAdam | 67.35 | 51.88 |
| 69.00(+1.65) | 56.03(+4.15) | |
| FedAvgM | 61.50 | 31.31 |
| 62.97(+1.47) | 33.13(+1.82) | |
| FedAdagrad | 69.33 | 49.22 |
| 71.17(+1.84) | 49.56(+0.34) |
| Prompt | Reg. | Hateful-Memes | CrisisMMD |
| 69.93 | 53.45 | ||
| ✓ | 73.51(+3.58) | 53.86(+0.41) | |
| ✓ | 73.41(+3.48) | 53.82(+0.37) | |
| ✓ | ✓ | 73.69(+3.76) | 58.46(+5.01) |
6 Conclusion and Discussion
Conclusion. In this work, we presented a multimodal heterogeneity benchmark to evaluate the performance of federated fine-tuning for MLLMs across five FL baselines on two downstream datasets, incorporating more than ten multimodal heterogeneities across four multimodal scenarios. To tackle the challenges posed by complex multimodal heterogeneities, we have also designed two simple yet effective modality-agnostic strategies. In the proposed FedMLLM framework, we conduct a series of experiments to highlight the advantages of FL over local training, examine the impact of multimodal heterogeneities, and evaluate the effectiveness of the two modality-agnostic strategies.
Discussion. FedMLLM currently focuses on the most fundamental and essential image-text multimodal data. However, other modalities—such as video and audio—still require further exploration in future work. When introducing video modalities, especially long videos, a key challenge is achieving efficient training and deployment given the limited computational resources on client devices. Additionally, the methods used in this benchmark to construct different modality scenarios are relatively simple. We hope that future benchmark studies will incorporate more realistic data distributions to better reflect real-world scenarios.
Supplementary Material
Gradient Norm and Loss Curves. FedAvg [27] and FedProx [22] methods are used to train the FedMLLM framework. However, due to the inability of average aggregation to effectively handle the impact of heterogeneity, LoRA experiences unstable training, causing both the loss and gradient norm to gradually increase, thereby hindering the effective convergence of the global model. In Fig. 8 and Fig. 9, the gradient norms and loss curves for both FedAvg and FedProx show a gradual upward trend. While the regularization term in FedProx helps partially mitigate this increase, it ultimately falls short in addressing the challenges posed by heterogeneity in the training of FedMLLM.
More Visualization of Heterogeneity. In the main text, we show the visualizations of the heterogeneity of aligned modal scenarios for the CrisisMMD dataset and the multimodal heterogeneity visualizations for the four multimodal scenarios in the Hateful-Memes dataset, as shown in Fig. 3 and Fig. 4. Furthermore, Fig. 10 and Fig. 11 supplement the visualizations of aligned modal heterogeneity and provide additional multimodal heterogeneity visualizations for the four multimodal scenarios in the Hateful-Memes dataset. Since the multimodal heterogeneity in the CrisisMMD dataset is similar to that in the Hateful-Memes dataset, redundant visualizations are omitted here for brevity.
As shown in Fig. 3 and Fig. 10, each dataset in our benchmark includes three levels of data heterogeneity. As the value decreases, the differences in data categories and quantities among clients increase. The visualizations of multimodal heterogeneities are presented in Fig. 4 and Fig. 11. In the aligned modal scenario, each client contains samples with all modalities. Based on data heterogeneity with , we further construct scenarios with multimodal heterogeneity under different non-aligned modal conditions. In the missing modal scenario, varying missing rates result in different proportions of missing modalities in the samples, although each client’s dataset still contains samples with all modalities. In the cross modal scenario, the proportion of single-modal clients varies. In the hybrid modal scenario, clients consist of both single-modal and multimodal combinations.
Complete Formatting Instructions. To prepare the fine-tuning data, the data format must be adjusted according to the prompt template shown in Tab. 9. Each sample should be structured as a dictionary containing the keys: “id”, “image”, and “conversations”. The “image” key specifies the path to the image, while “conversations” includes two dictionaries: one representing the user’s instruction, which contains the question and related information, and the other representing the assistant’s (MLLM) response. Since the task primarily involves classification, the response consists of the option letter and the corresponding category. For more details on data construction, please refer to the provided code.